# A Survey on Data Selection for Language Models

### v2

@@ 4 Data Selection for Instruction-Tuning and Multitask Training
Instruction-tuning and multitask training are methods for addressing the mismatch between the pretraining objective and a downstream user’s objective. Multitask training is a method where the model is trained on a wide variety of supervised tasks with the goal of performing all the training tasks, and possibly generalizing to unseen tasks. Recently, instruction-tuning has become a dominant training paradigm where the model is trained on pairs of (Instruction, Output), where Instruction denotes a human instruction for the model, and Output is the desired output, or an example of a desirable output. The goal of instruction tuning is to allow the model to constrain the models outputs in a way that is more controllable and helpful for the user (Zhang et al., 2023c).
Both multitask training and instruction-tuning generally assume that the model has been pretrained and has already acquired fundamental language capabilities. The intent of models that have been trained with either multitask training or instruction-tuning is, broadly speaking, to handle a very wide variety of possible inputs for downstream use cases as either classification or generative models. Thus, the aim of data selection for these settings is almost always focused on collecting a wider variety of data, and diversifying the data that already exists.
@@
As model sizes have scaled up and model capacity has correspondingly improved, distribution diversification has proven to be an effective method of improving model generalization, for both specialized tasks and general purpose model performance (Wei et al., 2021). Methods of multitask training (e.g. MT-DNN (Liu et al., 2019), Muppet (Aghajanyan et al., 2021), T5 (Raffel et al., 2020), ExT5 (Aribandi et al., 2021)) demonstrated the benefits of training on a range of datasets and tasks, selected for their perceived high-quality curation. Subsequently, a number of works (Khashabi et al., 2020; McCann et al., 2018; Keskar et al., 2019) proposed unifying the format of any language task into questions and answers, effectively adjusting the data point characteristics, with the goal of better matching the train and test distributions and expanding the number of tasks a model could learn in tandem, further improving model generalization. Based on this premise, Natural Instructions (Mishra et al., 2021), FLAN 2021 (Wei et al., 2021), and P3 (Sanh et al., 2022) scaled up the number and diversity of training tasks, distinguished by templated instructions—now known as instruction tuning.
At this stage, data selection was based mainly on availability across open-sourced dataset repositories (e.g. Hugging Face and TensorFlow Datasets), and pressure to scale, rather than any principled selection criteria. To extend this scaling beyond the openly available data, data augmentation techniques such as translation (Muennighoff et al., 2022b; Yong et al., 2022; Dhole et al., 2021; Singh et al., 2024; Üstün et al., 2024), input inversion (Min et al., 2022; Longpre et al., 2023a), and negative examples (Wang et al., 2022b) have successfully been used to further enrich the data variety. A number of works also increase data diversity by templatizing data in multiple prompt formats, including zero-shot, few-shot, and chain-of-thought prompts (Chung et al., 2022; Iyer et al., 2022). These works demonstrate that task variety and template variety both drive greater generalization, though Wei et al. (2021) report easier training tasks, such as sentiment analysis, contribute less.
@@
More recently, instruction-tuning data has been further diversified through the use of existing models using conversation as the unified data format. Training data can be generated directly by a model, often from the OpenAI API, creating examples in a loop (Honovich et al., 2022; Wang et al., 2023c; Xu et al., 2023; Taori et al., 2023; Lee et al., 2023; Kim et al., 2023)Wang et al. (2023c) begin with a small seed pool of example tasks, which the model uses to iteratively generate new tasks and instances of that task, followed by filters to ensure that data is “high-quality” and to remove duplicated data. Evol-instruct methods enhance these techniques to iteratively re-write training instances to be more complex and challenging (Xu et al., 2023). Lu et al. (2023a) iteratively trains on and produces its own training data with self-refinement techniques. These methods are shown to produce a diverse array of prompts that can be guided to match the distribution of certain real world tasks, domains, and topics which are not covered in existing datasets such as P3, FLAN, or the dataset used to train OPT-IML.
A survey of synthetic data tasks showed that they tend to focus on longer-form responses than what exists in academic datasets, as well as tasks related to brainstorming, planning, explanation, and creativity (Longpre et al., 2023b). The advent of synthetic data has also spurred developers to filter out system-specific outputs, such as disclaimers and refusals (e.g. “I’m sorry, but[…]” [^15] https://huggingface.co/datasets/teknium/GPT4-LLM-Cleaned from OpenAI models).
@@ 5 Data Selection for Preference Fine-tuning: Alignment
Various alignment methods, referred to under the umbrella of Reinforcement Learning from Human Feedback (RLHF), RL from AI Feedback (RLAIF), or Direct Preference Optimization (DPO) methods, involve the integration of human preferences into model behavior. This training process is designed to steer model responses to be generally more helpful and less harmful, whereas in other training phases such as pretraining or instruction-tuning, these preference signals may not be clearly defined within the utility function. These methods, grouped under Preference Fine-tuning (PreFT), typically follow instruction-tuning in the training pipeline of large generative models. The format of this data is often trios of (Prompt; Chosen, Rejected), where a Prompt is an instruction or other request from a user, Chosen is the preferred answer and Rejected is the lesser answer. Some methods work on groups of responses to a prompt, but they are less prevalent, and recently, Ethayarajh et al. (2024) proposed Kahneman-Tversky Optimization (KTO) which requires only the chosen or rejected continuation, but not both. The benefit of PreFT methods is to remove the requirement of strict next-token or accuracy-based predictions from ML models and integrate other, sometimes nuanced, qualitative signals instead.
RLHF and RLAIF methods have a central step which is the independent training of a reward model for active selection or weighting of samples during training. A reward model is a separate model fine-tuned to take in any input, such as text, and output a scalar reward indicating the utility (as represented by the probability that the text would be chosen). Data selection methods can be applied via training a reward model or the downstream policy, except for Direct Preference Optimization (DPO) methods that directly optimize the reward model and policy jointly from one set of data (Rafailov et al., 2023).
Data selection methods for PreFT are very nascent and are often focused on obtaining signal on specific capabilities and evaluations from the model. In this setting, the primary approaches to data selection are manual filtering, model-based evaluation, and reward model re-weighting (e.g. rejection sampling). The simplicity of the data selection methods reflects the relative underspecification of the goals of preference training, where basic questions are actively being studied on how best to collect preference data, and for what purpose (Lambert et al., 2023a; Casper et al., 2023). This area is also studied in Liu et al. (2023b) primarily for aligning instruction-tuned models, where the authors focus on three axes: complexity, quality, and diversity.
@@
Initial methods for collecting and filtering data for PreFT involved manually selecting data and using heuristic utility functions. For example, LIMA (Zhou et al., 2023) utilized post score, response length, formatting, topic, and other heuristics to manually select content from StackExchange, wikiHow, and Reddit to formulate 1000 high-quality training datasets for superficial alignment. Multiple works have indicated the use of external organizations to procure their data (Touvron et al., 2023b; Nakano et al., 2021; Bai et al., 2022b; Rajani et al., 2023), where detailed instructions are written by researchers and sent to a data-labeling company. However, it is common practice for data-labeling companies to filter the data internally before sending it to the model training organization and these processes are not documented beyond general notions of controlling for diversity and quality. Furthermore, this process is not reproducible. Ethayarajh et al. (2022) filter Reddit data based on the number of votes received on comments and posts in the comment chain when creating the Stanford Human Preferences (SHP) dataset. Similarly, Bai et al. (2022a) and Lambert et al. (2023b) weigh data points based on a measure of minimum engagement from users when creating their respective datasets. UltraFeedback, a popular dataset used to train state-of-the-art chat models such as Zephyr-$\beta$ (Tunstall et al., 2023) and Tulu 2 (Ivison et al., 2023b), implements a multi-aspect filtering process to create a preference dataset focused on instruction following, helpfulness, truthfulness, and honesty (Cui et al., 2023a). Further manual filtering has been performed on this data, and others such as Orca DPO pairs [^16] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs to improve the performance by verifying accuracy (Bartolome et al., 2023). The authors perform stratified sampling across a variety of existing preference datasets, filter for length, and decontamintate their data. Additionally, they use a set of 17 models to obtain diversity in their dataset.
@@
While human oversight can be higher quality, the ability of models to accurately label undesirable (e.g. incorrect) data has improved significantly, leading to an increased usage of language models to assign utility. For example, the Phi series of models (Gunasekar et al., 2023) use filtering to achieve “textbook quality” data from the web, as performed by a custom LM-based classifier, combined with synthetic data generated from OpenAI models. UltraFeedback (Cui et al., 2023a) and Nectar (Zhu et al., 2023) use GPT-4 to filter and/or rank the responses for quality. Critic models (Shepherd (Wang et al., 2023a) and Prometheus (Kim et al., 2023)) are used to critique model responses and identify diverse errors, and can be used to rank data points for quality. An example of using a model to filter data is by asking an LLM if an edited piece of text fixed a critique, and if the answer is no, the datapoint is removed, as done in Castricato et al. (2024). They are used as a part of the RLAIF cycle, where the critic is used as a reward model, to rank multiple model outputs. As this area of methods matures, it is likely that such critic models will form a central part of the data selection process by simply removing erroneous data.
@@
Emerging methods have applied reward-weighted data selection within a feedback loop, more akin to earlier forms of RLHF, such as with DPO models (Yuan et al., 2024b), or methods imitating self-play with less clear data selection methods (Chen et al., 2024). For example, Liu et al. (2024b) performs a rejection-sampling-like optimization with statistical weightings predicting the likelihood of whether a completion came from the optimal policy or base SFT policy to improve downstream preference optimization similar to DPO or Sequence Likelihood Calibration (SLiC). Especially in preference-based alignment, the border between synthetic data generation and data selection are often blurred as best practices are determined.
@@ 6 Data Selection for In-Context Learning
In-context learning (ICL) is a widely used prompting paradigm for language models (LMs). Instead of being fine-tuned with the LM, a few demonstration examples are given as a prompt to guide the language model to perform a similar prediction task on the input query (Brown et al., 2020). In-context learning is known to be sensitive to the choice (Perez et al., 2021), and even ordering (Lu et al., 2022), of the demonstrations. To improve the ICL performance without extensively training the potentially large LM, many recent papers have worked on constructing better in-context demonstrations by: selecting an optimal ordering from a fixed set of demonstrations, selecting from a large set of labeled data, or strategically annotating a small set of unlabeled data.
@@
Some works have focused on selecting a fixed set of demonstrations that achieves the best average performance across all testing inputs sampled from a target task. For example, Zhang et al. (2022) use reinforcement learning techniques to learn a policy that assigns utility to demonstrations for use on previously unseen tasks. A different approach, proposed by Wang et al. (2023b), is to use a smaller pre-trained language model to learn task-defined latent tokens and then select the examples that can best reconstruct the latent tokens.
@@
Many other works propose retrieval-based methods, where they train a demonstration retriever that can select demonstrations specifically tailored to each individual test data point. A straightforward method is retrieving the most similar examples to the testing input by cosine similarity of sentence embeddings (Liu et al., 2022). Gao et al. (2023b) further enhance this approach by retrieving demonstrations whose label lies in top-2 zero-shot predictions of the testing input. Rubin et al. (2022) score the similarity-based retrieved candidate demonstrations by their one-shot in-context learning performance for sequence-to-sequence tasks, then train a retriever with the collected scores. Li et al. (2023c) advance this framework through unified training across various datasets. Hashimoto et al. (2023) analyze the effect of different utility functions in training demonstration retrievers, and propose that incremental utility, which is the 1-shot utility score minus the 0-shot utility score, is better than directly using the 1-shot utility score. Ye et al. (2023) improve upon the previous works by scoring a set of demonstrations instead of only one demonstration. Finally, when prompting a model for semantic parsing, Levy et al. (2023) propose to select diverse demonstrations that collectively cover all of the structures required for the output program. For a comprehensive survey on this line of work, please see Xu et al. (2024b).
@@
There are also (iterative) sampling-based methods that do not require training a retriever to select the demonstrations. The most straightforward approach would be to uniformly sample a set of demonstrations from the candidate set as used by Min et al. (2022). Chang & Jia (2022) train a datamodel to predict the validation performance for a set of chosen demonstrations and select the demonstrations with the highest validation performance. Similarly, Nguyen & Wong (2023) propose constructing a validation set and evaluating each training data point by contrasting the validation performance of prompts with and without that data point, and formalizing the performance difference as influence. Li & Qiu (2023) iteratively select in-context demonstrations by an informativeness score, which is the language model prediction probability difference between one-shot and zero-shot inputs, and a diversity score, which discounts semantically similar examples. Gupta et al. (2023) propose to select the most informative set of demonstrations, instead of individual informative demonstrations, with a greedy algorithm to maximize the coverage of the chosen demonstrations. Similarly, Xu & Zhang (2024) iteratively refine the selected set of demonstrations by computing the misconfidence score, which is the ratio between the highest probability that the language model assigned to any incorrect label, and the output probability of the correct label.

### v3

@@ 4 Data Selection for Instruction-Tuning and Multitask Training
Instruction-tuning and multitask training are methods for addressing the mismatch between the pretraining data distribution and downstream use cases. Multitask training is a method where the model is trained on a wide variety of supervised tasks with the goal of performing all the training tasks and possibly generalizing to unseen tasks. Recently, instruction-tuning has become a dominant training paradigm where the model is trained on pairs of (Instruction, Output), where Instruction denotes a human instruction for the model, and Output is the desired output, or an example of a desirable output. The goal of instruction tuning is to encourage the model to generate outputs in a way that is more controllable and helpful for the user (Zhang et al., 2023c).
Both multitask training and instruction-tuning generally assume that the model has been pretrained and has already acquired fundamental language capabilities. The intent of models that have been trained with either multitask training or instruction-tuning is, broadly speaking, to handle a very wide variety of possible inputs for downstream use cases as either classification or generative models. Thus, the aim of data selection for these settings is typically focused on collecting a wider variety of data, and diversifying the data that already exists.
@@
As model sizes have scaled up and model capacity has correspondingly improved, distribution diversification has proven to be an effective method of improving model generalization, for both specialized tasks and general purpose model performance (Wei et al., 2021). Methods of multitask training (e.g., MT-DNN (Liu et al., 2019), Muppet (Aghajanyan et al., 2021), T5 (Raffel et al., 2020), ExT5 (Aribandi et al., 2021)) demonstrated the benefits of training on a range of datasets and tasks, selected for their perceived high-quality curation. Subsequently, a number of works (Khashabi et al., 2020; McCann et al., 2018; Keskar et al., 2019) proposed unifying the format of any language task into questions and answers, effectively adjusting the data point characteristics, with the goal of better matching the train and test distributions and expanding the number of tasks a model could learn in tandem, further improving model generalization. Based on this premise, Natural Instructions (Mishra et al., 2021), FLAN 2021 (Wei et al., 2021), and P3 (Sanh et al., 2022) scaled up the number and diversity of training tasks, distinguished by templated instructions—now known as instruction tuning.
At this stage, data selection was based mainly on availability across open-sourced dataset repositories (e.g., Hugging Face and TensorFlow Datasets), and pressure to scale, rather than any principled selection criteria. To extend this scaling beyond the openly available data, data augmentation techniques such as translation (Muennighoff et al., 2022b; Yong et al., 2022; Dhole et al., 2021; Singh et al., 2024; Üstün et al., 2024), input inversion (Min et al., 2022; Longpre et al., 2023a), and negative examples (Wang et al., 2022b) have successfully been used to further enrich the data variety. A number of works also increase data diversity by templatizing data in multiple prompt formats, including zero-shot, few-shot, and chain-of-thought prompts (Chung et al., 2022; Iyer et al., 2022). These works demonstrate that task variety and template variety both drive greater generalization, though Wei et al. (2021) report easier training tasks, such as sentiment analysis, contribute less.
@@
More recently, instruction-tuning data has been further diversified through the use of existing models using conversation as the unified data format. Training data can be generated directly by a model, often from the OpenAI API, creating and expanding examples iteratively (Honovich et al., 2022; Wang et al., 2023c; Xu et al., 2023; Taori et al., 2023; Lee et al., 2023; Kim et al., 2023)Wang et al. (2023c) begin with a small seed pool of example tasks, which the model uses to iteratively generate new tasks and instances of that task, followed by filters to ensure that data is “high-quality” and to remove duplicated data. Evol-instruct methods enhance these techniques to iteratively re-write training instances to be more complex and challenging (Xu et al., 2023). Lu et al. (2023a) iteratively trains on and produces its own training data with self-refinement techniques. These methods are shown to produce a diverse array of prompts that can be guided to match the distribution of certain real world tasks, domains, and topics which are not covered in existing datasets such as P3, FLAN, or the dataset used to train OPT-IML.
A survey of synthetic data tasks showed that they tend to focus on longer-form responses than what exists in academic datasets, as well as tasks related to brainstorming, planning, explanation, and creativity (Longpre et al., 2023b). The advent of synthetic data has also spurred developers to filter out system-specific outputs, such as disclaimers and refusals (e.g., “I’m sorry, but[…]” [^16] https://huggingface.co/datasets/teknium/GPT4-LLM-Cleaned from OpenAI models).
@@ 5 Data Selection for Preference Fine-tuning: Alignment
Various alignment methods, which include Reinforcement Learning from Human Feedback (RLHF), RL from AI Feedback (RLAIF), and Direct Preference Optimization (DPO), involve the integration of human preferences into model behavior. This training process is designed to steer model responses to be generally more helpful and less harmful, whereas in other training phases such as pretraining or instruction-tuning, these preference signals may not be clearly defined within the utility function. These methods, grouped under Preference Fine-tuning (PreFT), typically follow instruction-tuning in the training pipeline of large generative models. The format of this data is often trios of (Prompt; Chosen response, Rejected response), where a Prompt is an instruction or other request from a user, Chosen is the preferred answer and Rejected is the less-preferred answer. Some methods work on groups of responses to a prompt, but they are less prevalent, and recently, Ethayarajh et al. (2024) proposed Kahneman-Tversky Optimization (KTO) which requires only the chosen or rejected continuation, but not both. The benefit of PreFT methods is to remove the requirement of strict next-token or accuracy-based predictions from ML models and integrate other, sometimes nuanced, qualitative signals instead.
RLHF and RLAIF methods include the important step of independent training of a reward model for active selection or weighting of samples during training. A reward model is a separate model fine-tuned to take in any input, such as text, and output a scalar reward indicating the utility (as represented by the probability that the text would be chosen). Data selection methods can be applied via training a reward model or the downstream policy, except for Direct Preference Optimization (DPO) methods that directly optimize the reward model and policy jointly from one set of data (Rafailov et al., 2023).
Data selection methods for PreFT are very nascent and are often focused on obtaining signal on specific capabilities and evaluations from the model. In this setting, the primary approaches to data selection are manual filtering, model-based evaluation, and reward model re-weighting (e.g., rejection sampling). The simplicity of the data selection methods reflects the relative underspecification of the goals of preference training, where basic questions are actively being studied on how best to collect preference data, and for what purpose (Lambert et al., 2023a; Casper et al., 2023). This area is also studied in Liu et al. (2023b) primarily for aligning instruction-tuned models, where the authors focus on three axes: complexity, quality, and diversity.
@@
Initial methods for collecting and filtering data for PreFT involved manually selecting data and using heuristic utility functions. Multiple works have indicated the use of external organizations to procure their data (Touvron et al., 2023b; Nakano et al., 2021; Bai et al., 2022b; Rajani et al., 2023), where detailed instructions are written by researchers and sent to a data-labeling company. However, it is common practice for data-labeling companies to filter the data internally before sending it to the model training organization and these processes are not documented beyond general notions of controlling for diversity and quality. Furthermore, this process is not reproducible. Ethayarajh et al. (2022) filter Reddit data based on the number of votes received on comments and posts in the comment chain when creating the Stanford Human Preferences (SHP) dataset. Similarly, Bai et al. (2022a) and Lambert et al. (2023b) weigh data points based on a measure of minimum engagement from users when creating their respective datasets. UltraFeedback, a popular dataset used to train state-of-the-art chat models such as Zephyr-$\beta$ (Tunstall et al., 2023) and Tulu 2 (Ivison et al., 2023b), implements a multi-aspect filtering process to create a preference dataset focused on instruction following, helpfulness, truthfulness, and honesty (Cui et al., 2023a). Further manual filtering has been performed on this data, and others such as Orca DPO pairs [^17] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs improve the performance by verifying accuracy (Bartolome et al., 2023). The authors perform stratified sampling across a variety of existing preference datasets, filter for length, and decontamintate their data. Additionally, they use a set of 17 models to obtain diversity in their dataset.
@@
While human oversight can be higher quality, the ability of models to accurately label undesirable (e.g., incorrect) data has improved significantly, leading to an increased usage of language models to assign utility. For example, the Phi series of models (Gunasekar et al., 2023) use filtering to achieve “textbook quality” data from the web, as performed by a custom LM-based classifier, combined with synthetic data generated from OpenAI models. UltraFeedback (Cui et al., 2023a) and Nectar (Zhu et al., 2023) use GPT-4 to filter and/or rank the responses for quality. Critic models (Shepherd (Wang et al., 2023a) and Prometheus (Kim et al., 2023)) are used to critique model responses and identify diverse errors, and can be used to rank data points for quality. An example of using a model to filter data is by asking an LLM if an edited piece of text fixed a critique, and if the answer is no, the datapoint is removed, as done in Castricato et al. (2024). They are used as a part of the RLAIF cycle, where the critic is used as a reward model, to rank multiple model outputs. As this area of methods matures, it is likely that such critic models will form a central part of the data selection process by simply removing erroneous data.
@@
Emerging methods have applied reward-weighted data selection within a feedback loop, more akin to earlier forms of RLHF, such as with DPO models (Yuan et al., 2024b), or methods imitating self-play with less clear data selection methods (Chen et al., 2024). For example, Liu et al. (2024b) performs a rejection-sampling-like optimization with statistical weightings predicting the likelihood of whether a completion came from the optimal policy or base SFT policy to improve downstream preference optimization similar to DPO or Sequence Likelihood Calibration (SLiC) (Zhao et al., 2022). Especially in preference-based alignment, the border between synthetic data generation and data selection are often blurred as best practices are determined.
@@ 6 Data Selection for In-Context Learning
In-context learning (ICL) is a widely used prompting paradigm for language models (LMs). Instead of being used to fine-tune the LM, a few demonstration examples are given as a prompt to guide the language model to perform a similar prediction task on the input query (Brown et al., 2020a). In-context learning is known to be sensitive to the choice (Perez et al., 2021), and even ordering (Lu et al., 2022), of the demonstrations. To improve the ICL performance without training the potentially large LM, many recent papers have worked on constructing better in-context demonstrations by: selecting an optimal ordering from a fixed set of demonstrations, selecting from a large set of labeled data, or strategically annotating a small set of unlabeled data.
@@
Some works have focused on selecting a fixed set of demonstrations that achieves the best average performance across all testing inputs sampled from a target task. For example, Zhang et al. (2022a) use reinforcement learning techniques to learn a policy that assigns utility to demonstrations for use on previously unseen tasks. A different approach, proposed by Wang et al. (2023b), is to use a smaller pre-trained language model to learn task-defined latent tokens and then select the examples that can best reconstruct the latent tokens.
@@
Many other works propose retrieval-based methods, where they train a demonstration retriever that can select demonstrations specifically tailored to each individual test data point. A straightforward method is retrieving the most similar examples to the testing input by cosine similarity of sentence embeddings (Liu et al., 2022). Gao et al. (2023b) further enhance this approach by retrieving demonstrations whose label lies in top-2 zero-shot predictions of the testing input. Rubin et al. (2022) score the similarity-based retrieved candidate demonstrations by their one-shot in-context learning performance for sequence-to-sequence tasks, then train a retriever with the collected scores. Li et al. (2023c) advance this framework through unified training across various datasets. Hashimoto et al. (2023) analyze the effect of different utility functions in training demonstration retrievers, and propose that incremental utility, which is the 1-shot utility score minus the 0-shot utility score, is better than directly using the 1-shot utility score. Additionally, Iter et al. (2023) show that the cross entropy difference between candidate demonstrations and the test example is a promising selection utility. Ye et al. (2023) improve upon the previous works by scoring a set of demonstrations instead of only one demonstration. Finally, when prompting a model for semantic parsing, Levy et al. (2023) propose to select diverse demonstrations that collectively cover all of the structures required for the output program. For a comprehensive survey on this line of work, please see Xu et al. (2024b).
@@
There are also (iterative) sampling-based methods that do not require training a retriever to select the demonstrations. The most straightforward approach would be to uniformly sample a set of demonstrations from the candidate set as used by Min et al. (2022). Chang & Jia (2022) train a datamodel (Ilyas et al., 2022) to predict the validation performance for a set of chosen demonstrations and select the demonstrations with the highest validation performance. Similarly, Nguyen & Wong (2023) propose constructing a validation set and evaluating each training data point by contrasting the validation performance of prompts with and without that data point, and formalizing the performance difference as influence. Li & Qiu (2023) iteratively select in-context demonstrations by an informativeness score, which is the language model prediction probability difference between one-shot and zero-shot inputs, and a diversity score, which discounts semantically similar examples. Gupta et al. (2023) propose to select the most informative set of demonstrations, instead of individual informative demonstrations, with a greedy algorithm to maximize the coverage of the chosen demonstrations. Similarly, Xu & Zhang (2024) iteratively refine the selected set of demonstrations by computing the misconfidence score, which is the ratio between the highest probability that the language model assigned to any incorrect label, and the output probability of the correct label.